Please help me with
June is 1 year younger than her husband Lee. The product of there ages is 650. How old is each?
x(x+1)=650

Found 2 solutions by scott8148, checkley71:
Answer by scott8148(6628)

you are looking for two factors of 650 that are 1 apart

65 and 10? no ... 50 and 13? no ... but half of 50 is 25 and twice 13 is 26 ...BINGO!

the ages are 25 and 26

looking for factors that are close together, sometimes the square root helps ... the square root of 650 is 25 and a fraction

Answer by checkley71(8403)

X(X+1)=650
X^2+X-650=0
(X+26)(X-25)=0
X-25=0
X=25 AGE OF JANE.
25+1-26 AGE OF LEE.